Grade 12Analytical Geometry

the st line makes equal angle to three mutually perpendicular st lines whose dcs are given how can i find its direction cosines

aarshabh
9 Years agoGrade 12

1 Answer

Nishant Vora
9 Years ago
LEt us take our coordinate system in such a way that coordinate axes lie on thosethree mutually perpendicular st lines, Then Dc of req line will be(1/\sqrt{3},1/\sqrt{3},1/\sqrt{3})